# Aspects of Particle Production from Bubble Dynamics at a First Order Phase Transition

### v2

@@ I Introduction
In addition to GWs, such phase transitions can also be responsible for particle production: any particle that couples to the background scalar field undergoing the phase transition gets produced due to the spacetime dynamics of the scalar field over the duration of the transition. Note that particle production should be a significantly stronger phenomenon than GW emission since the graviton coupling is Planck suppressed, whereas particle couplings tend to be significantly larger. Furthermore, while GWs can only be produced after the bubbles collide as they correspond to transverse, traceless excitations of the metric, which cannot be sourced by spherically symmetric bubbles, particle production is more general and can occur over all stages of FOPTs, including the nucleation and expansion of spherical bubbles. In the GW literature, particle production from FOPTs is primarily considered in the context of interactions between the fast moving bubble walls and the surrounding plasma [48, 49, 50, 51, 52, 53, 54]. While such plasma-induced particle production effects can be significant, in this paper we concern ourselves with the more fundamental, irreducible form of particle production that occurs purely due to the dynamics of the background field itself. Particle production from a changing background field is a well-known physical phenomenon familiar from various contexts, such as gravitational particle production [55, 56, 57, 58], Schwinger effect [59], or Hawking radiation from black holes [60, 61]. Such particle production, which exists even in the absence of a plasma, could be particularly relevant for phase transitions in vacuum, or supercooled transitions [62, 63, 64, 65, 66, 67, 68, 69, 70]. In cases where a thermal plasma is present, the details of particle production can still be important for deducing the GW spectra as well as for various non-thermal particle production applications to BSM physics, in particular since the bubble walls can be ultrarelativistic, producing particles with masses far heavier than the scale of the phase transition or the temperature of the thermal bath (as has been explored, e.g. in [71, 72, 73, 74, 75]).
@@ II.1 Homogeneous Case: Bogoliubov Transformation
It is important to keep in mind that these formulae are only applicable if the relevant particle gets its mass from the background field. For fields with bare masses much larger than that contributed by the background field, the production gets exponentially suppressed.
@@ II.2 Particle Production at Bubble Nucleation
with $\mathcal{I}(a)$ given by Eq. (8).
@@ IV.2 Simple Cases
Next, consider the case where two planar bubble walls, each with speed $v_{\mathrm{w}}$, approach each other and undergo a perfectly elastic collision, getting reflected with the same speeds $v_{w}$, re-establishing the false vacuum in the region in between (see Fig. 1(b), with the collision occurring at the origin). This corresponds to the perfectly elastic collision configuration studied in [76, 77, 71]. The field configuration is this case can be written as
@@
Figure 1: Spacetime evolution of some simple field configurations of planar thin-walled bubbles, plotted for wall speed $v_{w}\approx 1$. The yellow (light) and blue (dark) regions corresponds to the true and false vacua respectively.
@@
Incredibly, this is the same as the Fourier transform for the case of an elastic bubble collision, Eq. (23), with half the amplitude (originating from the fact that the scalar field only changes in half the plane in this case). This suggests that the mode excitations in Eq. (23) originate from the “turning around” of the wall at the origin. In other words, rather than some complicated dynamics from the collision of the two walls, the mode excitations arise from each wall decelerating and turning around at the intersection point.
@@
The simple field ansatz used in the above examples, in Eq. (22) and Eq. (24), which match the ones used in the earlier studies [76, 71], miss one crucial ingredient. For both, the field value evaluates to $v_{\phi}$ at the origin. However, when the two bubbles walls completely superpose, the field value is known to add linearly, thus getting excited away from the true minima, jumping up to $2v_{\phi}$ [99]. This piece is missing in the above treatments. We can incorporate this additional ingredient by adding to the field configuration a sharp jump at the origin, such as a sharp Gaussian of the form $e^{-(t^{2}+x^{2})/\sigma}$, where $\sigma\gg 1$ determines how sharply the Gaussian is peaked around the origin. Since the Fourier transform is linear, the overall result can be obtained simply by calculating the Fourier transform of the above Gaussian and adding it to the previous results, e.g. Eq. (23) and Eq. (25). Evaluating this numerically, we find that the Fourier transform of this Gaussian peak gives vanishing contributions to $\omega^{2}-k^{2}>0$ modes. This is also consistent with the numerical results from the realistic cases studied in [78] discussed below, which did consider field excitations to $2v_{\phi}$ at the collision point.
@@ VI Summary
Overall, particle production from the dynamics of the background scalar field is found to be a significant aspects of first order phase transitions yet remains relatively understudied, especially compared to the more broadly and extensively studied phenomenon of gravitational waves production. Such particle production holds important implications for several relevant aspects: it can affect the dynamics of the background field, provide an additional source of gravitational waves related to FOPTs (this will be explored in future work [108]), as well as address outstanding problems in particle physics and cosmology, such as dark matter and baryogenesis (as explored in [71, 72, 73, 74, 75]), and therefore remains a crucial ingredient of FOPTs that merits further attention. In this context, the improved understanding of the particle production process at various stages of a FOPT provided by this paper invites further investigation into various related matters. In particular, several results here, such as particle production from accelerating bubbles, the choice of the parameters in the case of non-universal masses and interactions, as well as the arguments for the inefficiency of nonperturbative processes, are based on simple intuitive arguments, but could (and should) be further sharpened with more rigorous calculations.
@@ Appendix A Particle Production from a Single Propagating Bubble Wall
$(g^{\alpha\beta}\partial_{\alpha}\partial_{\beta}-m^{2})\phi=0$ (39)
@@
$\left[\,\left(\frac{1-v_{w}^{2}}{1+v_{w}^{2}}\right)\left(-\frac{\partial^{2}}{\partial x_{2}^{2}}+\frac{\partial^{2}}{\partial x_{1}^{2}}\right)-\frac{4v_{w}}{1+v_{w}^{2}}\frac{\partial}{\partial x_{1}}\frac{\partial}{\partial x_{1}}\right]\phi(t,x)={m_{\psi}^{2}(x_{1})}\phi(t,x)\,,$ (40)
@@
The physics is now invariant under translations along $x_{2}$, hence its conjugate momentum $p_{2}$ is a conserved quantity. We can therefore expand the field $\psi$ in terms of its $p_{2}$ modes, and solve the modified KG equation Eq. 40 mode-by-mode (note that such momentum decomposition is only justified if the corresponding momentum is conserved, hence this choice of coordinates) to construct our solution. For each $p_{2}$ mode, Eq. 40 gives
$\left(\frac{\partial^{2}}{\partial x_{1}^{2}}-\frac{4iv_{w}p_{2}}{1-v_{w}^{2}}\frac{\partial}{\partial x_{1}}\right)\phi(x_{1})=\left[\frac{1+v_{w}^{2}}{1-v_{w}^{2}}{m(x_{1})^{2}}-p_{2}^{2}\right]\phi(x_{1})\,.$ (42)
@@
$e^{-i\omega t+ipx}=\text{exp}\left[-\frac{\omega v_{w}+p}{\sqrt{1+v_{w}^{2}}}x_{1}-\frac{\omega-pv_{w}}{\sqrt{1+v_{w}^{2}}}x_{2}\right]$ (43)

### v3

@@ I Introduction
In addition to GWs, such phase transitions can also be responsible for particle production: any particle that couples to the background scalar field undergoing the phase transition gets produced due to the spacetime dynamics of the scalar field over the duration of the transition. Note that particle production should be a significantly stronger phenomenon than GW emission since the graviton coupling is Planck suppressed, whereas particle couplings tend to be significantly larger. Furthermore, while GWs can only be produced after the bubbles collide as they correspond to transverse, traceless excitations of the metric, which cannot be sourced by spherically symmetric bubbles, particle production is more general and can occur over all stages of FOPTs, including the nucleation and expansion of spherical bubbles. In the GW literature, particle production from FOPTs is primarily considered in the context of interactions between the fast moving bubble walls and the surrounding plasma [48, 49, 50, 51, 52, 53, 54]. While such plasma-induced particle production effects can be significant, in this paper we concern ourselves with the more fundamental, irreducible form of particle production that occurs purely due to the dynamics of the background field itself. Particle production from a changing background field is a well-known physical phenomenon familiar from various contexts, such as gravitational particle production [55, 56, 57, 58], Schwinger effect [59], or Hawking radiation from black holes [60, 61]. Such particle production, which exists even in the absence of a plasma, could be particularly relevant for phase transitions where a plasma is absent (such as phase transitions in vacuum) or diluted (such as supercooled transitions (see e.g. [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73])). In cases where a thermal plasma is present, the details of particle production can still be important for deducing the GW spectra as well as for various non-thermal particle production applications to BSM physics, in particular since the bubble walls can be ultrarelativistic, producing particles with masses far heavier than the scale of the phase transition or the temperature of the thermal bath (as has been explored, e.g. in [74, 75, 76, 77, 78]).
@@ II.1 Homogeneous Case: Bogoliubov Transformation
The Bogoliubov transformation calculation assumes that the background change is sudden (i.e. non-adiabatic), otherwise the field simply tracks the change in the background field smoothly and remains in the vacuum state, so that no particle production occurs. The condition for non-adiabaticity is $|\omega^{\prime}/\omega^{2}|\geq 1$. Since the mass changes by $m$ over a timescale $\mathcal{T}$, we have $\omega^{\prime}\sim m/\mathcal{T}$, whereas $\omega^{2}\approx m^{2}$. Thus the evolution is non-adiabatic only for $m\mathcal{T}\leq 1$, and the results below are only valid in this regime.
@@
It is important to keep in mind that these formulae are only applicable if the relevant particle gets its mass from the background field. For fields with bare masses much larger than that contributed by the background field, the production gets exponentially suppressed; this occurs when the non-adiabaticity condition is violated, i.e. when $\delta m/m\ll m\mathcal{T}$.
@@ II.2 Particle Production at Bubble Nucleation
with $\mathcal{I}(a)$ given by Eq. (8). Again, this formula, derived from a Bogoliubov tranformation calculation, is only valid when the evolution is non-adiabatic, which occurs only for $l_{w}\,m_{X}\leq 1$.
@@ IV.2 Simple Cases
Next, consider the case where two planar bubble walls, each with speed $v_{\mathrm{w}}$, approach each other and undergo a perfectly elastic collision, getting reflected with the same speeds $v_{w}$, re-establishing the false vacuum in the region in between (see Fig. 1(b), with the collision occurring at the origin). This corresponds to the perfectly elastic collision configuration studied in [79, 80, 74]. The field configuration in this case can be written as
@@
Figure 1: Spacetime evolution of some simple field configurations of planar thin-walled bubbles, plotted for wall speed $v_{w}\approx 1$, with $t,x$ in units of $v_{\phi}^{-1}$. The yellow (light) and blue (dark) regions corresponds to the true and false vacua respectively.
@@
Incredibly, this is the same as the Fourier transform for the case of an elastic bubble collision, Eq. (23), with half the amplitude (originating from the fact that the scalar field only changes in half the plane in this case). This suggests that the mode excitations in Eq. (23) originate from the “turning around” of the wall at the origin. In other words, rather than some complicated dynamics from the collision of the two walls, the mode excitations arise from each wall decelerating and turning around at the intersection point. While this case is included here primarily for heuristic reasons, physically it corresponds to the case where a bubble nucleates instantaneously and expands out, and the result is consistent with particle production only at nucleation, with no particle production taking place during the expansion phase. The underlying physics of this result will be clarified in the final paragraph of this subsection (after case (d) below).
@@
The simple field ansatz used in the above examples, in Eq. (22) and Eq. (24), which match the ones used in the earlier studies [79, 74], miss one crucial ingredient. For both, the field value evaluates to $v_{\phi}$ at the origin. However, when the two bubbles walls completely superpose, the field value is known to add linearly, thus getting excited away from the true minima, jumping up to $2v_{\phi}$ [102]. This piece is missing in the above treatments. We can incorporate this additional ingredient by adding to the field configuration a sharp jump at the origin, such as a sharp Gaussian of the form $e^{-(t^{2}+x^{2})/\sigma}$, where $\sigma/v_{\phi}^{2}\gg 1$ determines how sharply the Gaussian is peaked around the origin. Since the Fourier transform is linear, the overall result can be obtained simply by calculating the Fourier transform of the above Gaussian and adding it to the previous results, e.g. Eq. (23) and Eq. (25). Evaluating this numerically, we find that the Fourier transform of this Gaussian peak gives vanishing contributions to $\omega^{2}-k^{2}>0$ modes. This is also consistent with the numerical results from the realistic cases studied in [81] discussed below, which did consider field excitations to $2v_{\phi}$ at the collision point.
@@ VI Summary
Overall, particle production from the dynamics of the background scalar field is found to be a significant aspects of first order phase transitions yet remains relatively understudied, especially compared to the more broadly and extensively studied phenomenon of gravitational waves production. Such particle production holds important implications for several relevant aspects: it can affect the dynamics of the background field, provide an additional source of gravitational waves related to FOPTs [111], as well as address outstanding problems in particle physics and cosmology, such as dark matter and baryogenesis (as explored in [74, 75, 76, 77, 78]), and therefore remains a crucial ingredient of FOPTs that merits further attention. In this context, the improved understanding of the particle production process at various stages of a FOPT provided by this paper invites further investigation into various related matters. In particular, several results here, such as particle production from accelerating bubbles, the choice of the parameters in the case of non-universal masses and interactions, as well as the arguments for the inefficiency of nonperturbative processes, are based on simple intuitive arguments, but could (and should) be further sharpened with more rigorous calculations.
@@ Appendix A Particle Production from a Single Propagating Bubble Wall
$(g^{\alpha\beta}\partial_{\alpha}\partial_{\beta}-m^{2})\psi=0$ (39)
@@
$\left[\,\left(\frac{1-v_{w}^{2}}{1+v_{w}^{2}}\right)\left(-\frac{\partial^{2}}{\partial x_{2}^{2}}+\frac{\partial^{2}}{\partial x_{1}^{2}}\right)-\frac{4v_{w}}{1+v_{w}^{2}}\frac{\partial}{\partial x_{1}}\frac{\partial}{\partial x_{2}}\right]\psi(x_{1},x_{2})={m_{\psi}^{2}(x_{1})}\psi(x_{1},x_{2})\,,$ (40)
@@
The physics is now invariant under translations along $x_{2}$, hence its conjugate momentum $p_{2}$ is a conserved quantity. We can therefore expand the field $\psi$ in terms of its $p_{2}$ modes, which are plane waves, and solve the modified KG equation Eq. 40 mode-by-mode (note that such momentum decomposition is only justified if the corresponding momentum is conserved, hence this choice of coordinates) to construct our solution. For each $p_{2}$ mode, Eq. 40 gives the following equation for its coefficient, which we write as $\psi_{p2}(x_{1})$:
$\left(\frac{\partial^{2}}{\partial x_{1}^{2}}-\frac{4iv_{w}p_{2}}{1-v_{w}^{2}}\frac{\partial}{\partial x_{1}}\right)\psi_{p2}(x_{1})=\left[\frac{1+v_{w}^{2}}{1-v_{w}^{2}}{m_{\psi}(x_{1})^{2}}-p_{2}^{2}\right]\psi_{p2}(x_{1})\,.$ (42)
@@
$e^{-i\omega t+ipx}=\text{exp}\left[-\frac{\omega v_{w}+p}{\sqrt{1+v_{w}^{2}}}x_{1}-\frac{\omega-pv_{w}}{\sqrt{1+v_{w}^{2}}}x_{2}\right]\,.$ (43)